A sum of $25,000 is divided among peter , queen and Raymond.
Raymond received half . Peter receives $3,125 and Queen receives the remainder. Calculate
The ratio in which the $25,000

Answer by math_tutor2020(3816)


Your question seems to be cut off. Though I could be mis-interpreting things.

I'll try to answer what I can.

The sum of $25,000 is being distributed to 3 people.
Raymond gets half, so he gets 25000/2 = 12,500 dollars.

Peter gets $3,125
So far this accounts for 12500+3125 = 15,625 dollars
And there's 25,000-15,625 = 9,375 dollars left over for Queen to get.

Raymond = $12,500
Peter = $3,125
Queen = $9,375

From here, it's not clear what your teacher wants you to do with this information. The instructions are missing. You'll need to please post the full problem, with full instructions. Include any relevant diagrams or tables if necessary.

Edit:
I realize now what your teacher is asking. My apologies
The ratio of the three amounts is 3125:9375:12500 when considering Peter, Queen and Raymond in that exact order. The ratio fully reduces to 1:3:4

Notice how the portions for Peter and Queen (1 and 3) add to 1+3 = 4, which is half the total parts of 1+3+4 = 8
Also, notice that 3125/25000 = 1/8